OpenStudy (anonymous):

Given the function f(x) = log2(x + 6), find the value of f^−1(3).

OpenStudy (anonymous):

I really need help with understanding this.

OpenStudy (anonymous):

Answer Choices: a. f−1(3) = 2 b. f−1(3) = 3 c. f−1(3) = 9 d. f−1(3) = 18

OpenStudy (anonymous):

@zepdrix help please

OpenStudy (anonymous):

@SkaterBoyShawn can you help me?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay good

OpenStudy (freckles):

\[f^{-1}(3)=a \implies f(a)=3 \\ \text{ so you need to solve } \log_2(a+6)=3 \text{ for } a \]

OpenStudy (anonymous):

@freckles in my notes it was saying how i had to find the inverse function or something like that

OpenStudy (anonymous):

let me solve it 1sec

OpenStudy (anonymous):

ok @SkaterBoyShawn

OpenStudy (freckles):

You can find the inverse function then plug in 3 or you can just solve the equation above for a

OpenStudy (anonymous):

ugh im still confused

OpenStudy (anonymous):

i think i got the answer

OpenStudy (freckles):

on what part?

OpenStudy (freckles):

like why \[f^{-1}(3)=a \implies f(a)=3 ? \\ \text{ or on solving } \log_2(a+6)=3 \text{ for } a?\]

OpenStudy (anonymous):

how? @SkaterBoyShawn and @freckles everything like what do you do with log2?

OpenStudy (freckles):

you want to write in the equivalent exponential form

OpenStudy (freckles):

recall \[\log_b(x)=y \implies b^{y}=x\]

OpenStudy (anonymous):

i dont get any of this at all. Math is my weakest subject

OpenStudy (freckles):

so you don't know how to compare \[\log_b(x)=y \text{ to } \log_2(a+6)=3 \\ \text{ then use that } \log_b(x)=y \implies b^y=x \\ \text{ to write } \log_2(a+6)=3 \text{ in exponential form }\]?

OpenStudy (freckles):

 what is in place of the b? in place of the x? in place of the y?